Feb 25, 2007
The meeting was called to order at approximately 18:00 UT. Those attending: Roger, (chairing President), Gerard, Steve, and Nancy.
- Agenda Item 1: Minutes of Last Meeting - Approval
- - Since there were no objections, the Minutes of the last Board Meeting were considered ratified as correct.
- Agenda Item 2: Annual Elections Congradulations
- - Re-elected Officers and Board members were congradulated
- Agenda Item 3: Projects Review Discussion
- - Roger reported orders are steady, and inventory holding, 500 CD's and 300 booklets in stock.
- - When inventory drops below 200 in either stock, the base will be reworked ahead of reprinting or burning additional CD's to include new
information and bring them current.
- - Nancy stated she intends to do another germination test on seed stock not tested last year. This is to be able to advise the public during
interviews on how to keep their personal stocks fresh. .
- - Roger reported that the server is currently idle, as he moved and has yet to reconnect. However, there is no demand for the server at
present. There is no problem with the server, however.
- - Nancy reported that her Blog radio shows continue weekly, but rebroadcasts from older shows are being done as her schedule has picked
up. She will be on TV and doing live radio interviews.
- Agenda Item 4: Adjournment
- - The meeting was adjourned at 18:39 UT.